dy/dx=x+(y/(x^2)) where y(1)=1.5 and x=1(0.3)1.6 advanced euler method

dy/dx=x+(y/(x^2))

where

y(1)=1.5

and

x=1(0.3)1.6

advanced euler method
Answer #1

solution: y O=16 x² Given that dy = nt y In we need to find y(1.6) with step size h: 0:3 Here finiy) = nt y n² No = 1, yo = 12 2nd appr zoximation : Ti = y + hf (my, you 2.232204 2.23.2204 + (0-3) 113 + 3 ) (1:3)2 Now = 2.232204 +0.3) (2.620831) 8 =